A whale swimming against an ocean current traveled 72 mi in 3 h. Swimming in the opposite direction, with the current, the whale was able to travel the same distance in 1.8 h.
Find the speed of the whale in calm water and the rate of the ocean current.

Answer by checkley77(12844)

D=RT
R=D/T
R=72/3
R=24 MPH. AGAINST THE CURRENT.
R=72/1.8
R=40 MPH. WITH THE CURRENT.
(40-24)/2=16/2=8 MPH. IS THE CURRENT.
PROOF:
40-8=32 MPH. IS THE SWIMMING SPEED OF THE WHALE IN CALM WATERS.
24+8=32 MPH. DITTO